The length, X, in minutes of a MS Teams video lecture for STBE101 is normally distributed


with a mean of 28 minutes and standard deviation of 2 minutes.


(a) Find 𝑃𝑃(20 < 𝑋𝑋 < 28)


(b) Find 𝑃𝑃(𝑋𝑋 > 33)

1
Expert's answer
2021-11-12T16:11:35-0500

(a)


"P(20<X<28)=P(X<28)-P(X\leq 20)"

"=P(Z<\dfrac{28-\mu}{\sigma})-P(Z\leq \dfrac{20-\mu}{\sigma})"

"=P(Z<\dfrac{28-28}{2})-P(Z\leq \dfrac{20-28}{2})"

"=P(Z<0)-P(Z\leq -4)"

"=0.5-0.00003167"

"=0.49996833"

(b)


"P(X>33)=1-P(X\leq 33)"

"=1-P(Z\leq \dfrac{33-\mu}{\sigma})"

"=1-P(Z\leq \dfrac{33-28}{2})"

"=1-P(Z\leq2.5)"

"=0.00620967"